Axolotl:

This is an urgent math question because I have limited time! Please help :) Please show steps on how you go the answer f(x) = −16x2 + 60x + 16 Part A: What are the x-intercepts of the graph of f(x)? Show your work.

Axolotl:

This is urgent because I have maybe less than an hour rn

aWaY:

ok do 16 x 2 first then add that by 16 then whatever u get plug it in with 60x

aWaY:

i think thats how but i could be wrong

aWaY:

did you get it?

Axolotl:

for some reason my brain does not like this math

Axolotl:

and it is 16^2 sorry not 16x2

Axolotl:

16x^2

Vocaloid:

To solve for the x intercepts, you need to find the value(s) where f(x) = 0 f(x) = −16x2 + 60x + 16 Factoring out -4 = -4(4x^2 - 15x - 4) Factoring again (use the cross method) = -4(x - 4)(4x + 1) = 0 From there, set (x-4) = 0 and (4x + 1) = 0 and solve for x